OpenStudy (anonymous):

the base of a triangle is 5cm greater than the height the area is 18 cm ^2 find the height and length of the base

OpenStudy (anonymous):

(x)(x+5)=2*18 x=4

OpenStudy (anonymous):

x(x+5)=36 x^2+5x-36=0 (x+9)(x-4)=0 x=4 so base 4hight 9

OpenStudy (anonymous):

dipankar do you mean height 4cm length 9cm?

OpenStudy (anonymous):

yes...